SUBMILLIMETER-WAVE SPECTRUM OF CH_{2}D^{+}

\underline{T. AMANO}, Department of Chemistry and Department of Physics and Astronomy, University of Waterloo, Waterloo, ON, Canada N2L 3G1.

In interstellar carbon chemistry, CH_{3}^{+} is thought to be an important and abundant molecular ion. However, as it is a symmetric planar molecule and, as a result, it has no permanent dipole moment, it is almost impossible to detect this species by radio astronomical observations. Its deuterated species, CH_{2}D^{+} and CHD_{2}^{+}, possess the dipole moment, so the rotational lines should be observable. Rösslein et al.^{a} and Jagod et al.^{b} observed the infrared spectra of these deuterated species. Demuynck and coworkers^{c} tried to observe CH_{2}D^{+} rotational lines in an extended negative glow discharge with no success. More recently Lis et al.^{d} reported tentative identification of CH_{2}D^{+} toward Ori IRc2. The molecular constants and the predicted rotational transition frequencies given by Röslein et al.^{a} were a good starting point in searching for the rotational lines. A very weak feature was found almost exactly at the calculated frequency for the 2_{12} −1_{11} transition. Eventually the line appeared stronger enough for precise frequency measurements, after adjusting the reaction conditions. The optimum gas mixture was found to be CH_{4} (∼ 3 mTorr), CD_{4} (∼ 1 mTorr), H_{2} (∼ 2 mTorr), and He (∼ 35 mTorr). It is interesting to note that helium is essential to produce CH_{2}D^{+}. No signals were detectable with Ar buffer. Although the signal was seen without H_{2}, it appears to play a subtle role in the formation, resulting in about a factor 2 increase in intensity. Adding D_{2} instead of CD_{4} resulted in no signal. The observations were made with about 16 mA discharge current with liquid nitrogen cooling. As this ion is a light molecule and the signal was only weakly observed, four transitions were detected so far in the 280-890 GHz region. All observed transition frequencies agree within 1MHz of the predicted frequencies. These laboratory transition frequencies strongly support the tentative astronomical identification by Lis et al.^{d}

aM. Rösslein et al., Astrophys. J. 382, L51 (1991)

bM.-F. Jagod et al., J. Mol. Spectrosc. 153, 666 (1992)

cC. Demuynck, J. Mol. Spectrosc. 168, 215 (1994)

dD. C. Lis et al., in Submillimeter Astrophysics and Technology, ASP Conference Series, 417, 23 (2009)